Out of 800 families with 4 children each, about how many of these families would have at least one boy?
Answer by mathmate(429)

Question:
Out of 800 families with 4 children each, about how many of these families would have at least one boy?

Solution:
The probability of a 4-child family having at least one boy is equal to
P(n>=1)=1-P(n=0)=1-(1/2)^4=15/16.
Out of 800 families, the most likely number is 800*15/16=750 families with at least one boy.